OpenStudy (anonymous):

HELP PLEASE ILL GIVE MEDAL AND FAN!!!! XD Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (anonymous):

@susanaM09

OpenStudy (anonymous):

0,1,1,2,3

OpenStudy (anonymous):

why? can you explain?

OpenStudy (anonymous):

f(1),f(2),f(3),f(4),f(5)

OpenStudy (anonymous):

okay that doesn't make much sense to me

OpenStudy (anonymous):

you have f(1) and f(2) f(3)=f(2)+f(1) f(4)=f(3)+f(4) f(5)=f(4)+f(3)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thanks for helping me out !!!

OpenStudy (anonymous):

yw